# RealityCraft: An In-Situ CAD+CAM Interface for Novices via Scene-Aware Augmented Reality

### v1

@@ Abstract.
Despite the growing accessibility of augmented reality (AR) for visualization, existing computer-aided design systems remain largely confined to traditional screens and are often inaccessible to novice users due to their complexity. We present RealityCraft, an open-sourced AR interface that enables in-situ computer-aided design and manufacturing (CAD+CAM) for novices. Unlike traditional CAD systems confined to computer screens, RealityCraft allows users to design directly within their physical environments, with primitive geometries. RealityCraft recognizes and utilizes physical constraints such as furniture and walls, enhancing user interaction through spatial awareness and depth occlusion. Furthermore, RealityCraft features an integrated AR-based 3D printing workflow, where users can drag and drop designs onto their 3D printer’s virtual twin in their immediate space. Through a user study, we demonstrate that RealityCraft enhances engagement and ease of use for novices. By bridging the gap between digital creation and physical output, RealityCraft aims to transform everyday spaces into creative studios.
@@ 1. Introduction
The introduction of Tinkercad AR111https://www.tinkercad.com/ipad-app, Accessed: 16.07.2024 marked a notable progression by incorporating AR to view designs in physical spaces, albeit mainly for visualization. Applications such as Gravity Sketch222https://www.gravitysketch.com/, Accessed: 16.07.2024 have begun to leverage virtual reality (VR) headsets and controllers to enable more immersive 3D sketching experiences. In recent HCI research, several notable advancements have emerged to enhance spatial and intuitive design interfaces. pARam (Stemasov et al., 2024) suggests that in-situ parametric design and customization can be effectively achieved through hand interactions and environmental considerations in extended reality (XR). Additionally, BrickStARt (Stemasov et al., 2023a) introduces a novel approach by integrating tangible construction blocks with mixed reality (MR) to enable users to design and test physical artifacts directly in their intended locations. ProtoAR (Nebeling et al., 2018) addresses the challenges of AR content creation by combining physical prototyping with mobile authoring tools for rapid development and interaction. Furthermore, MixFab (Weichel et al., 2014) illustrates how MR can streamline personal fabrication by facilitating direct interaction with virtual objects within an AR environment. Collectively, these advancements highlight a trend towards more spatially aware and intuitive design interfaces, which are particularly beneficial for complex tasks that integrate digital and physical spaces.
@@
Despite these advancements, a significant gap remains in harnessing AR for interactive design that is fully integrated with the user’s physical environment and intuitive to use. Focusing on this gap, we introduce RealityCraft, an innovative AR interface that enables in-situ CAD+CAM for novices. Our system bridges the gap between simple, novice-friendly tools and the immersive, spatial capabilities of advanced AR technologies. RealityCraft uses intuitive, constructive solid geometry modeling with primitives similar to those in Tinkercad, but enhances the design experience by integrating these elements within the user’s physical environment through AR. This approach makes the design process more intuitive and also adaptable to the physical constraints and possibilities of the environment.
@@
Our system utilizes a state-of-the-art headset (Meta Quest 3333https://www.meta.com/quest/quest-3/, Accessed: 16.07.2024) and its hand-gesture controls as well as passthrough features, enabling users to interact with and create 3D models directly atop their actual desks or on other flat surfaces like walls or the ground (Figure 1a). This setup not only facilitates natural interaction with the digital content but also allows users to accurately gauge the size and scale of their designs in relation to their immediate surroundings. Such features are critical in creating a truly immersive and intuitive design experience that blends the digital with the physical.
Unlike other tools, RealityCraft is designed to recognize and adapt to real-world scene elements like furniture, enhancing its utility and applicability in everyday spaces, letting users turn any surface into their CAD workspace. It incorporates depth occlusion and scene awareness to ensure that digital objects realistically interact with the physical world, which is a notable advancement over other AR or VR systems that lack this level of integration.
@@
Further, RealityCraft extends the concept of AR in CAD beyond mere visualization to include actual manufacturing processes. It integrates seamlessly with 3D printing technology, incorporating a drag-and-drop metaphor that allows users to transition effortlessly from digital design to physical creation (Figure 1c-e).
@@
To validate the effectiveness of RealityCraft, we conducted a user study (n=7) with three design tasks. Our findings indicate that RealityCraft enhances user engagement and supports novice users in design tasks. The results suggest high usability, with users reporting ease of use and a low learning curve.
@@
Overall, we present a novel design system that integrates spatial and in-situ features with intuitive hand gesture controls and a straightforward design logic tailored for novice users. Our contributions can be summarized as follows:
We introduce RealityCraft, a novel in-situ CAD+CAM interface for novices that enables 3D design directly within users’ physical environments through hand-gesture interactions and real-time scene recognition.
We present a set of spatial interaction techniques for modeling and fabrication, allowing users to manipulate 3D primitives and work within context-aware physical spaces, incorporating depth occlusion and furniture identification for improved immersion, realism, and accuracy.
We integrate 3D printing workflows directly into the AR design process, enabling a seamless transition from digital design to physical creation through a drag-and-drop interface with virtual twin representations of 3D printers.
@@
We conduct a user study evaluating the effectiveness of RealityCraft, demonstrating high user engagement and performance, especially for novice users with no prior CAD experience.
@@
We make RealityCraft open-source444The link was removed for anonymity., contributing to the research community by providing the system and codebase for further development, experimentation, and possible integration with additional fabrication technologies.
@@
Our approach simplifies geometric modeling and supports a flexible methodology for designing in any place, smoothly transitioning from digital designs to physical outputs. By enhancing accessibility and creativity through a context-aware and user-friendly platform, we aim to advance the field of design and fabrication through AR. Our system invites further exploration and innovation, which we hope paves the way for a future where physical and digital realms seamlessly converge.
@@ 2. Related Work
This section reviews the literature and previous work related to our study in three main areas: Extended Reality in digital design and fabrication, in-situ tools and interfaces, and systems tailored for novice users.
@@ 2.1. XR in Digital Design and Fabrication
Extended Reality technologies, which include Virtual Reality, Augmented Reality, and Mixed Reality, have increasingly been applied to enhance various aspects of digital design and fabrication.
@@
VR tools have been mostly developed for digital design compared to fabrication. Tools like Gravity Sketch allow users to create 3D models in an immersive environment, providing a spatial understanding that traditional screen-based interfaces cannot offer. VRSketchIn by Drey et al. (Drey et al., 2020) combines traditional 2D input methods (pen and tablet) with 3D VR sketching, leveraging the strengths of both to enhance precision and versatility. These systems serve as interim sketching tools that can facilitate the design workflow.
@@
AR has been explored across a broader range of systems in design and fabrication. While head-mounted displays (HMDs) can be used with hand gestures or controllers in AR (Guo and Prabhakaran, 2022), most AR systems rely on tablets with pen or finger interactions. For example, massless555https://massless.dev/, Accessed: 16.07.2024, focuses on reducing material waste during prototyping by overlaying design information directly onto the deployment area of the artifact, enhancing precision and minimizing errors due to physical constraints. Another example is Window-Shaping by Huo et al. (Huo et al., 2017), which enables rapid 3D model creation on and around physical objects in AR, facilitating design ideation without the need for 3D reconstruction or fiducial markers. Sketched Reality, proposed by Kaimoto et al. (Kaimoto et al., 2022), introduces an AR sketching tool that interacts with physically actuated robots, exploring the physics of designs and mechanisms. RealitySketch by Suzuki et al. (Suzuki et al., 2020) introduces a dynamic and responsive interface for sketching interactive graphics and visualizations. Other AR tools focus on prototyping models within their intended spaces (Wang et al., 2023; Alghofaili et al., 2023; Leiva et al., 2021, 2020), streamlining digital design and fabrication workflows.
@@
The integration of XR in digital design and fabrication also extends to MR environments. James and Eckert (James and Eckert, 2023) have developed MR systems that allow operators to visualize potential outcomes projected onto the material via real-time physics simulation, enabling adjustments before the actual cutting process begins. There are other works in the area of XR explored in spatially conveying information to users (Chen et al., 2023; Cheng et al., 2023; Dogan et al., 2024; Li et al., 2024; Han et al., 2023; Cheng et al., 2021; Lindlbauer et al., 2019). These can be integrated with digital design and fabrication systems to provide instructions or context-aware interactable menus.
The overlap between XR tools and digital fabrication has been discussed by Dogan et al. (Dogan et al., 2022) with the focal point of both being used for creating objects in an instant. This can be correlated to XR tools used for rapid prototyping including VRceptioin by Gruenefeld et al. (Gruenefeld et al., 2022), ProtoAR by Nebeling et al. (Nebeling et al., 2018), SpatialProto by Müller et al. (Müller et al., 2021), and VirtualComponent by Kim et al. (Kim et al., 2019), bringing design and fabrication closer to a more interactive form (Willis et al., 2010).
@@ 2.2. In-Situ Tools and Interfaces
In-situ tools and interfaces refer to systems designed to be used directly within the environment where a task or activity is taking place, rather than in a separate or abstract setting. Specifically, the in-situ design allows the creation or modification of designs within the context of their intended deployment. Recent advancements in AR technology have expanded the possibilities for in-situ design by allowing designers to visualize changes in the actual environment where the design will exist. This can eliminate the discrepancy between the site of artifact creation and its eventual placement in traditional digital design tools (Ashbrook et al., 2016; Mahapatra et al., 2019). There are many in-situ design tools with different unique features. DesignAR by Reipschläger and Dachselt (Reipschläger and Dachselt, 2019) combines interactive 2D screens with stereoscopic AR for refined object tracing. RoMA by Peng et al. (Peng et al., 2018) synchronizes design and fabrication, with the capability to incorporate existing objects into the workspace.
@@
The concept of in-situ is not only about overlaying information but also about interacting with the environment in a meaningful way. Teachable Reality by Monteiro et al. (Monteiro et al., 2023) lets users define their own in-situ tangible and gestural interactions in real time, allowing users to prototype functional AR applications without programming. HoloDesk by Hilliges et al. (Hilliges et al., 2012) and MixFab by Weichel et al. (Weichel et al., 2014) introduces low-friction immersive design tools with direct gestural controls and the ability to add physical objects into designs, making the design process more context-aware. Lau et al. introduced Modeling In Context (Lau et al., 2010), which uses everyday objects as a base point for designing around them. pARam by Stemasov et al. (Stemasov et al., 2024) provides in-situ parametric design through hand interactions. FusePrint by Zhu et al. (Zhu et al., 2016) explored design for fitting mounts, while Robiot by Li et al. (Li et al., 2019) introduces designing of mechanisms that actuate other objects, and Protopiper by Agrawal et al. (Agrawal et al., 2015) discussed room scale object design with low fidelity. Mitterberger et al. (Mitterberger et al., 2020; Mitterberger, 2022; Mitterberger et al., 2023) focus on larger-scale designs in architecture.
@@
Tools incorporating tangible artifacts with in-situ design are useful for context awareness and experimenting with physics, reducing the need for simulation. Stemasov et al. presented BrickStARt (Stemasov et al., 2023a), a low-fidelity tangible design tool using interlocking building blocks, providing an accessible way of prototyping for users. Letter and Wolf developed Tangible Version Control (Letter and Wolf, 2022), which explores versioning of physical artifacts with the help of an HMD. Printy3D by Yung et al. (Yung et al., 2018) targets the education of children in 3D modeling with tangible artifacts placed in a virtual container. Other tools include linkage (Jeong et al., 2018) or carving (Hattab and Taubin, 2019a, b), which makes the process interactive by touch while designing.
In-situ tools have been developed for consultation with makers. For instance, created for school students, RobotAR by Villanueva et al. (Villanueva et al., 2021) facilitated remote learning and collaboration through an AR-compatible robot, and WireDraw by Yue et al. (Yue et al., 2017) explored guiding a user’s manual fabrication using a 3D-printing pen with instructions delivered via an HMD. Other examples cover the potential for training workers, mainly by providing tools to subject matter experts for creating interactive training content (Chidambaram et al., 2021; Ipsita et al., 2021).

### v2

# TinkerXR: In-Situ, Reality-Aware CAD and 3D Printing Interface for Novices

@@ Abstract.
Despite the growing accessibility of augmented reality (AR) for visualization, existing computer-aided design (CAD) systems remain confined to traditional screens or require complex setups or predefined parameters, limiting immersion and accessibility for novices. We present TinkerXR, an open-sourced interface enabling in-situ design and fabrication through Constructive Solid Geometry (CSG) modeling. TinkerXR operates solely with a headset and 3D printer, allowing users to design directly in and for their physical environments. By leveraging spatial awareness, depth occlusion, recognition of physical constraints, reference objects, and intuitive hand movement controls, TinkerXR enhances realism, precision, and ease of use. Its AR-based workflow integrates design and 3D printing with a drag-and-drop interface for a 3D printer’s virtual twin. A user study comparing TinkerXR with TinkerCAD demonstrates higher accessibility, engagement, and ease of use for novices. By bridging the gap between digital creation and physical output, TinkerXR transforms everyday spaces into accessible and expressive creative studios.
@@ 1. Introduction
The introduction of Tinkercad AR111https://www.tinkercad.com/ipad-app, Accessed: 16.07.2024 marked a notable progression by incorporating AR to view designs in physical spaces, albeit mainly for visualization. Applications such as Gravity Sketch222https://www.gravitysketch.com/, Accessed: 16.07.2024 have begun to leverage virtual reality (VR) headsets and controllers to enable more immersive 3D sketching experiences. In recent HCI research, several notable advancements have emerged to enhance spatial and intuitive design interfaces. pARam (Stemasov et al., 2024) suggests that in-situ parametric design and customization can be effectively achieved through hand interactions and environmental considerations in AR. Additionally, BrickStARt (Stemasov et al., 2023a) introduces a novel approach by integrating tangible construction blocks with mixed reality (MR) to enable users to design and test physical artifacts directly in their intended locations. ProtoAR (Nebeling et al., 2018) addresses the challenges of AR content creation by combining physical prototyping with mobile authoring tools for rapid development and interaction. Furthermore, MixFab (Weichel et al., 2014) illustrates how MR can streamline personal fabrication by facilitating direct interaction with virtual objects within an AR environment. Collectively, these advancements highlight a trend towards more spatially aware and intuitive design interfaces, which are particularly beneficial for complex tasks that integrate digital and physical spaces.
@@
Despite these advancements, a significant gap remains in harnessing AR for interactive design that directly integrates with the user’s physical environment, balances expressivity and intuitiveness, supports seamless transitions from 3D modeling to 3D printing, and delivers all these capabilities in one cohesive package. Focusing on this gap, we introduce TinkerXR, an innovative AR interface that enables in-situ CAD and 3D printing for novices. Our system bridges the gap between simple, novice-friendly tools and immersive, spatial capabilities of advanced AR technologies, offering advanced spatial interaction techniques for designing 3D models directly in the physical environment using only hand controls. TinkerXR employs intuitive, Constructive Solid Geometry modeling with primitives similar to those in Tinkercad, but enhances the design experience by integrating these elements within the user’s physical environment through AR, making the design process more intuitive and adaptable to physical constraints and possibilities, while seamlessly integrating a 3D printing workflow.
@@
Our system utilizes a state-of-the-art headset (Meta Quest 3333https://www.meta.com/quest/quest-3/, Accessed: 16.07.2024) and its hand-gesture controls as well as passthrough features, enabling users to interact with and create 3D models directly atop their actual desks or on other flat surfaces like walls or the ground (Figure 3a). This setup not only facilitates intuitive interaction with the digital content but also allows users to accurately gauge the size and scale of their designs in relation to their immediate surroundings, enabling them to design in a way that considers and adapts to the space they’re working within.
Unlike other tools, TinkerXR is designed to recognize and adapt to real-world scene elements like furniture, enhancing its utility and applicability in everyday spaces, letting users turn any surface into their CAD workspace. Additionally, by designing around these setups, users can leverage features like snapping to grid and referencing smaller real-world objects, allowing users to design around them, making it easier to align and create models with precision, further enhancing the design process within their physical environment. Such features are critical in creating a truly immersive and intuitive design experience that seamlessly blends digital with physical.
It also incorporates depth occlusion, a technique that ensures digital objects are visually blocked by real-world objects based on their physical location and perspective, creating a more realistic and immersive experience. This is a notable advancement over other AR or VR systems, which may not achieve the same level of integration between the digital and physical worlds.
@@
Furthermore, TinkerXR extends the concept of AR in CAD beyond digital output to include actual manufacturing processes. It integrates seamlessly with 3D printing, incorporating a drag-and-drop metaphor that allows users to intuitively prepare models for printing (Figure 3c-e). Acting as a built-in slicing interface in AR, TinkerXR prepares models for printing while providing real-time feedback through a virtual 3D printer twin, allowing users to make adjustments and visualize the output. Users can also initiate the 3D printing operation directly from the system, ensuring a seamless transition from digital design to physical creation.
@@
To validate the effectiveness of TinkerXR, we conducted a user study (n=10) with three design tasks. In addition to TinkerXR, we compared the results with Tinkercad as ground truth, using the same three tasks to assess performance across both systems. The study also included a survey to evaluate usability, user engagement, and the learning curve. Our findings indicate that TinkerXR enhances user engagement and supports novice users in design tasks. The results suggest high usability, with users reporting ease of use and a low learning curve.
@@
Overall, we present a novel design and fabrication system that integrates spatial and in-situ features with intuitive hand control schemes and a straightforward design logic tailored for novice users. By bridging the gap between the digital and physical worlds, our system enables seamless design-to-manufacturing workflows. Our contributions can be summarized as follows:
We present TinkerXR, an innovative in-situ CAD system that enables hand-based CSG modeling directly in any physical environment, tailored for novices. The system’s novel contributions include interaction methods specifically designed for CSG design in real-world settings, such as snap-to-grid, designing on top of reference objects, and intuitive controls for applying traditional CAD functionalities with hand interactions.
We integrate 3D printing workflows directly into the AR design process, allowing users to seamlessly transition from digital design to physical creation. The system includes a built-in slicer, real-time feedback via a virtual 3D printer twin, and the ability to initiate printing directly from the interface, offering a cohesive design-to-manufacturing pipeline.
@@
We conduct a user study evaluating the effectiveness of TinkerXR when compared to Tinkercad, demonstrating high user engagement and immersiveness for novice users with no to little prior CAD experience.
@@
We make TinkerXR open-source444The link was removed for anonymity., contributing to the research community by providing the system and codebase for further development, experimentation, and possible integration with additional fabrication technologies.
@@
Our approach simplifies geometric modeling in AR and supports a flexible methodology for designing in any place, smoothly transitioning from digital designs to physical outputs. By enhancing 3D creativity and accessibility through a context-aware immersive platform, we hope to advance the field of design and fabrication through AR.
@@ 2. Related Work
This section reviews the literature in three main areas: Extended reality (XR) in digital design and fabrication, in-situ tools and interfaces, and systems tailored for novice users. We provide a comparative overview of existing AR-based CAD systems, highlighting our system’s unique capabilities as summarized in Table 1.
| Situated Modeling (Lau et al., 2012) | MixFab (Weichel et al., 2014) | RoMA (Peng et al., 2018) | DesignAR (Reipschläger and Dachselt, 2019) | pARam (Stemasov et al., 2024) | WindowShaping (Huo et al., 2017) | TinkerXR | |||||||||||||
✓ wireframe printing |
✓ | tablet | ✓ | ||||||||||||
not provided | not provided | not provided | not provided | not provided | not provided | ✓ |
@@ 2.1. XR in Digital Design and Fabrication
XR technologies, which include virtual reality (VR), augmented reality (AR), and mixed reality (MR), have increasingly been applied to enhance various aspects of digital design and fabrication (Dogan et al., 2022a).
@@
VR tools have been mostly developed for digital design compared to fabrication. Tools like Gravity Sketch allow users to create 3D models in an immersive environment, providing a spatial understanding that traditional screen-based interfaces cannot offer. Systems such as VRSketchIn (Drey et al., 2020), AdaptiBrush (Rosales et al., 2021), and HandPainter (Jiang et al., 2021) serve as interim sketching tools that can facilitate the design workflow. Researchers also studied the effectiveness of CAD tools in VR (Bourdot et al., 2010; Feeman et al., 2018; Cordeiro et al., 2019). While these tools provide immersive spatial understanding for 3D modeling, they are typically limited to virtual-only environments, separating the design process from the physical context of deployment.
@@
In contrast to VR, AR operates directly in physical environment, allowing users to see their designs in their intended real-world context (Iyer et al., 2025; Dogangun et al., 2024; Dogan et al., 2023c). While head-mounted displays (HMDs) can be used with hand gestures (as seen in TinkerXR) or controllers in AR (Guo and Prabhakaran, 2022; Stemasov et al., 2024; Peng et al., 2018; Jasche and Ludwig, 2020), most of the previous AR systems rely on tablets with pen or finger interactions. For instance, massless555https://massless.dev/, Accessed: 19.01.2025, focuses on reducing material waste during prototyping by overlaying design information directly onto the artifact’s deployment area, enhancing precision and minimizing errors due to physical constraints. Sketched Reality (Kaimoto et al., 2022) introduces an AR sketching tool that interacts with physically actuated robots, exploring the physics of designs and mechanisms. RealitySketch (Suzuki et al., 2020) introduces a dynamic and responsive interface for sketching interactive graphics and visualizations.
TinkerXR contributes to the vision of works such as (Wang et al., 2023; Alghofaili et al., 2023; Leiva et al., 2021, 2020), which streamline digital design and fabrication workflows. However, unlike these systems, TinkerXR uniquely integrates a 3D printing workflow directly within its AR interface, providing an all-in-one solution that eliminates the need for external slicing tools or connector software while supporting precise and fully-functional designs.
@@
James and Eckert (James and Eckert, 2023) developed MR systems that allow operators to visualize potential outcomes projected onto the material via real-time physics simulation, enabling adjustments before a metal-cutting process begins. Other works explored conveying information spatially to users (Chen et al., 2023; Cheng et al., 2023; Li et al., 2024; Han et al., 2023). These can be integrated with digital design and fabrication systems to provide instructions or context-aware interactable menus (Dogan et al., 2024; Cheng et al., 2021; Lindlbauer et al., 2019) Window-Shaping (Huo et al., 2017) enables rapid 3D model creation on and around physical objects using a sketch-and-inflate scheme using a tablet.
Building on this foundation, TinkerXR uniquely and efficiently integrates in-situ constructive solid geometry (CSG) modeling in a seamless 3D printing workflow within an immersive AR interface. Unlike existing systems, TinkerXR allows direct 3D design and printing in any workspace without requiring additional tools or post-processing software.
@@ 2.2. In-Situ Tools and Interfaces
In-situ tools refer to systems used directly within the environment where a task takes place, rather than in separate or abstract settings. Recent advancements in AR have enabled designers to visualize changes within the actual environment, bridging the gap between artifact creation and its intended deployment (Ashbrook et al., 2016; Mahapatra et al., 2019). For example, DesignAR (Reipschläger and Dachselt, 2019) combines 2D tablet screens with stereoscopic AR and a pen for precise object tracing. RoMA synchronizes design and fabrication using a 3D printing robotic arm that incorporates existing objects into the workspace. HoloDesk (Hilliges et al., 2012) and MixFab (Weichel et al., 2014) introduce immersive tools with gestural controls tracked by Kinect, enabling context-aware designs. Unlike systems requiring fixed setups or specialized hardware, TinkerXR allows design in and for any environment using only an off-the-shelf HMD. This portability makes TinkerXR a versatile tool for diverse settings, whether at home or outside.
@@
The concept of in-situ is not only about overlaying information but also about interacting with the environment in a meaningful way. Teachable Reality (Monteiro et al., 2023) lets users define their own in-situ tangible and gestural real-time interactions, allowing them to prototype functional AR applications without programming. SnapToReality (Nuernberger et al., 2016) enables real-time alignment of virtual objects to real-world constraints, such as edges and planar surfaces. FusePrint (Zhu et al., 2016) explored design for fitting mounts, while Robiot (Li et al., 2019) introduces designing of mechanisms that actuate other objects, and Protopiper (Agrawal et al., 2015) discusses room-scale design with low fidelity. Mitterberger et al. (Mitterberger et al., 2020; Mitterberger, 2022; Mitterberger et al., 2023) focus on larger-scale architectural designs.
Some in-situ modeling systems follow a design in and for anywhere ideology. Lau et al. introduced Modeling in Context (Lau et al., 2010), which uses everyday objects as a base point for designing around them. pARam (Stemasov et al., 2024) provides in-situ parametric design, while Situated Modeling (Lau et al., 2012) utilizes fiducial markers for robust object tracking. While these markers can impact scene aesthetics and unobtrusiveness, researchers showed it is also possible to hide them using specialized imaging techniques (Dogan et al., 2022b, 2023b). TinkerXR builds on this design in and for anywhere ideology but takes a novel direction by CSG modeling. Unlike sketching or parametric approaches, CSG modeling requires the integration of precise features such as snapping to a grid, object alignment, and robust control logic systems to enable functional and flexible design workflows. The adaptation of these traditionally desktop-bound features into an in-situ AR context is what allows TinkerXR to advance this ideology.
@@
Tools incorporating tangible artifacts with in-situ design are valuable for context awareness and experimenting with physics, reducing the need for simulation. BrickStARt used interlocking building blocks for this purpose (Stemasov et al., 2023a). Tangible Version Control (Letter and Wolf, 2022) explores versioning of physical artifacts with the help of an HMD. Printy3D (Yung et al., 2018) teaches 3D modeling with tangible artifacts placed in a virtual container. Other tools include linkage (Jeong et al., 2018) or carving (Hattab and Taubin, 2019a, b), which make the design process interactive by touch.
Our system integrates tangibility into the design experience in several ways. It incorporates reference object types, i.e., representations of object that serve as design references, equipped with functionalities to design around them. It also features depth occlusion for physical objects, providing a realistic interaction between the real and digital worlds.